OpenStudy (anonymous):

how to factorize this question?? (x+y)^2+(x+y)-20

OpenStudy (anonymous):

[(x+y)+5][(x+y)-4]

OpenStudy (anonymous):

heyalbie can do that becasue suppose i let u=x+y, then i have\[u^2+u-20\]which factors to \[=(u+5)(u-4)\]now just replace u with x+y to get\[=(x+y+5)(x+y-4)\]

OpenStudy (anonymous):

Now you don't have to make the substitution; I just did that to help you see what heyalbie did; if the sub helps than by all means use it, otherwise just write x+y whereever you would normally wrote a single letter